# How LLMs Learn to Reason: A Complex Network Perspective

### v2

@@ 1 Introduction
Figure 2: Visualizing the Structural Evolution from “Skill Islands" to the Concept Web. These network snapshots provide a direct, qualitative narrative of the concept web’s formation, corresponding to the different phases of training. In each subfigure, green, red, and blue dots indicate the question, answer, and intermediate nodes of the CoNet, respectively. The color of each directed edge represents the transition probability from the head to the tail and is drawn according to the colorbar in the rightmost column. Here, only edges with transition probabilities greater than $0.95$ are shown. (a) In the early fast-learning phase (Step 20), the model discovers a few short, disjointed reasoning paths, representing the first nascent “skill islands”. (b) At the onset of slow learning (Step 50), the system has proliferated into a maximal collection of disconnected islands. (c) Deep in the slow-learning phase (Step 800), these previously separate islands have coalesced into a single, giant connected component, forming the unified and expansive concept web.
2 The V-Shaped Signature as Evidence for a Sparse Concept Web
The V-shaped evolution of the response length is a well-demonstrated, quantifiable hallmark of deep structural reorganization during RLVR training (He et al., 2025; Luo et al., 2025a). To anchor our investigation, we first reproduce this signature by applying RLVR to the DeepSeek-R1-Distill-Qwen-1.5B model following the DeepScaleR protocol (Luo et al., 2025a). As shown in Fig. 1, our experiment recovers the characteristic V-shaped curve (Fig. 1).
To dissect the mechanism underlying this signature—a task intractable within the LLM itself—we deploy an extended, multi-task variant of the CoNet as a “computational microscope” (See Appendix D for details). CoNet instantiates the implicit reasoning graph discussed in the Introduction as a concrete, tractable model: it maps the LLM “semantic states” to abstract nodes in a fixed random graph and LLM “logical transitions” to learnable, probabilistic edges. This minimal model reduces the learning process to a tractable graph-traversal problem and, remarkably, reproduces the V-shaped curve observed in our LLM experiment (Fig. 1), validating it as a faithful proxy for uncovering the underlying mechanism.
(a) Correct Response Histogram
@@ 2 The V-Shaped Signature as Evidence for a Sparse Concept Web
Figure 3: A Sparse-Web Structure Necessitates Longer Reasoning Chains. This figure links the emergent sparse topology of the concept web to the observed increase in response length. (b, c) The color and marker scheme is identical to that used in Fig. 2, and only edges with transition probability greater than $0.95$ are retained. The largest connected component (a local region of which is shown here) remains sparse with average degree around 22 even as it grows from step $50$ to $800$. (a) Consequently, the distribution of solution path lengths, shown in both raw counts (histogram) and probability density (inset), shifts decisively to the right during this period, confirming that navigating the sparse backbone requires longer reasoning chains.
Armed with this confirmation, we now demonstrate that the V-shaped evolution of the response length arises from a two-stage process. The decreases of the response length in the initial fast-learning stage is dominated by local optimization. Given the random initialization of transition probabilities, shorter paths inherently have a higher probability of being sampled. Consequently, the model rapidly discovers short, efficient solutions to individual problems, giving rise to a proliferation of disconnected “skill islands” (as shown in Fig. 2(a)). This burst is quantified by the number of distinct solution clusters, which peaks right before the onset of slow learning (Fig. 2(b)). Stage one thus constitutes a parallel search for local efficiency that naturally shortens the average response length.
The subsequent slow-learning phase, where response length steadily increases, marks a fundamental shift to global integration. The primary task is no longer discovering new islands but weaving them into a single, expansive concept web (see Fig. 2(c)), defined to be the largest connected component in which every edge has a transition probability above $0.95$. As the concept web relentlessly grows, a critical structural property emerges: the web remains persistently sparse, with its average degree stabilizing near two (as shown in Fig. 3(c)). This provides direct, quantitative validation for our central Hypothesis (Hypothesis 1).
This emergent sparsity is the topological key to the V-shape’s rising slope. A network with such a low average degree is dominated by chain and tree-like structures, inherently lacking the shortcuts and redundant pathways found in denser graphs. Consequently, connecting two previously distant concepts requires traversing long, intermediate paths. The histograms in Fig. 3(a) provide clear evidence of this effect: As the sparse web expands, the distribution of solution lengths undergoes a decisive and sustained rightward shift, confirming that navigation on this sparse backbone necessitates longer reasoning chains. Thus, the V-shaped curve furnishes a robust macroscopic signature of an underlying topological shift: from parallel, local optimization on disconnected components to serial, global navigation across an expanding, sparse backbone.
Yet can this simple, largely tree-like model capture the full complexity of LLM reasoning? After all, an LLM’s autoregressive memory enables sophisticated behaviors such as self-correction or reflection (e.g., “Wait, let me rethink that…”)—that would manifest as cycles in the reasoning graph. We argue they are fully compatible(Minegishi et al., 2025). These reasoning cycles are not the dense, ubiquitous shortcuts of a highly connected graph. Instead, they are likely the cognitive manifestation of sparse, local loops—occasional, high-cost revisions on a much larger, fundamentally chain-like scaffold. The existence of a limited number of such cycles does not significantly alter the global average degree of the web, which remains close to two. With the sparse-web model established, this link between a macroscopic dynamic and a microscopic topology forms the foundation for the analysis that follows.
(a) CoNet
(b) LLM
(c) Step 400 (Before SFT)
(d) Step 400 (After SFT)
(e) Step 450 (Continued RLVR)
Figure 4: The Fragility of a Sparse Web: Catastrophic Forgetting and Fast Recovery. This figure demonstrates a key prediction of our sparse-web hypothesis: that the concept web is fragile, relying on critical bridge-like connections. (c-e) The color and shape conventions follow those of Fig. 2, and only a local view of the concept web is shown. Orange crosses on the edges indicate decreases in transition weights, with thicker crosses denoting more substantial decreases. The microscopic view in CoNet shows how a lightweight supervised fine-tuning (SFT) on a converged web (c) severs these critical bridges, causing the structure to fragment (d). Subsequent RLVR rapidly repairs these links (e). (a, b) This microscopic severing manifests as macroscopic catastrophic forgetting in both CoNet and the 1.5B LLM, where performance plummets upon initiating SFT. The subsequent fast recovery once RLVR resumes highlights the localized nature of the damage, underscoring the web’s structural fragility.
3 The Double-Edged Sword of Sparsity: from catastrophic Forgetting to Simulated Annealing scheme
From the graph theoretic point of view, the sparsity in the conceptual network implies low redundancy, fragility, and critical vulnerability. It also reflects a structural hypothesis for a well-known operational challenge: why models that have formed a concept web via RLVR are so susceptible to catastrophic forgetting when subsequently fine-tuned with supervised learning (Li et al., 2024). Our theory posits that this is not a widespread erasure of knowledge, but a specific topological failure: the surgical severing of critical, bridge-like connections.
To test this mechanistic hypothesis, we subjected models in the deep slow-learning phase to a lightweight in-distribution SFT (see Appendix H and I). As expected, the intervention triggered a sharp drop in overall performance in both CoNet and the 1.5B LLM (Fig. 44(a), 4(b)). Having reproduced this known macroscopic phenomenon, we leveraged CoNet’s transparency to inspect the microscopic aftermath.
The result is a direct validation of our structural explanation. The microscopic view reveals that the SFT catastrophic forgetting was not a global degradation of the web. Instead, SFT’s aggressive optimization of a few paths overwrote the policy at critical branching nodes, precisely breaking the single connection points for vast downstream subgraphs. Fig. 44(c)4(d) provides a stark visualization of this: a unified web is instantly fragmented, offering a clear, structural cause for the catastrophic performance drop.
However, this is where the sword’s second edge is revealed. The damage, while profound in its effect, is topologically localized. SFT does not erase the knowledge encoded in the now-disconnected subgraphs; it merely renders them unreachable. This localization explains the remarkable speed of recovery, which can be observed in Fig. 44(a)4(b). When RLVR training resumes, it triggers a quick rebound because the task is simply to “re-solder” a few broken connections to restore access to the latent scaffold as shown in Fig. 44(e).
This two-step process of localized disruption followed by rapid repair is strongly analogous to simulated annealing Kirkpatrick et al. (1983). The brief SFT phase acts as a high-temperature “heating” shock, breaking connections and allowing the system to escape a potentially suboptimal state. The subsequent RLVR phase then acts as a controlled “cooling”, guiding the system as it re-forms connections, ideally settling into a more robust configuration (Lai et al., 2025). This insight—that a known vulnerability can be repurposed as a deliberate annealing step—forms the foundation for the Annealed-RLVR algorithm, which we introduce in Sec. 5.
4 Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition
Although the overall reward curve in RLVR training looks smooth, our sparse-web theory shows that much more is happening underneath. As shown in the individual trajectories in Fig. 5, the network’s sparse structure creates two competing effects: (i) a forgetting process caused by frustration, which is strongest at the start of the slow-learning phase, and (ii) an opposite, phase-transition-like learning process that gradually builds and integrates knowledge during slow learning.
(a) CoNet
(b) LLM
Figure 5: Microscopic Mechanisms of RLVR: Forgetting by Frustration, Learning by Phase Transition The learning trajectories for individual problems in CoNet (a) and the 1.5B LLM (b) reveal a fundamental duality. (i) Frustration-Induced Forgetting: At the onset of slow learning [see inset in (a)], the intense competition for connections on a sparse web manifests as volatile, non-monotonic accuracy curves, where some skills are competitively suppressed. (ii) Phase-Transition-Like Learning: Subsequently, the web’s sparse frontier enables new skills to be acquired in punctuated, accelerated jumps (orange and green). The smoother gradients in the LLM (vs. the clean steps in CoNet) can be attributed to finite-size effects, as explained in the main text.
4.1 Frustration-Induced Forgetting: The Price of Integration
The first dynamic is driven by frustration-induced forgetting: a competitive process where the advancement of some skills forces the deterministic regression of others as the model integrates its “skill islands”. This process manifests as volatile, non-monotonic learning curves in the learning dynamics (Fig. 5) and represents an internal “push-and-pull” fundamentally different from the external shock of SFT-induced catastrophic forgetting (Shenfeld et al., 2025; Jin et al., 2025a).
This process peaks in the maximally frustrated state at the onset of the slow-learning stage, a critical juncture where “skill islands” compete for limited connections under the $\left<k\right>\approx 2$ constraint. Paradoxically, this state maximizes exploration: the model’s ability to find diverse, correct out-of-distribution solutions peaks, far exceeding the base model or the same model after prolonged training (Fig. 6(b), blue). This subsequent degradation in performance is consistent with the phenomenon of policy collapse, which has been recently reported in several studies to occur after extended RLVR, leading to a loss of exploratory power (Chen et al., 2025; Yue et al., 2025; Cui et al., 2025).
4.2 Phase-Transition-Like Learning: The Engine of Growth
The opposing dynamic, phase-transition-like learning, is an important engine of new knowledge acquisition (Power et al., 2022; Lin et al., 2025). The evidence for this is clearly shown in CoNet’s sharp, step-like accuracy jumps (Fig. 5(a)), and the steep accelerated learning gradients observed in the 1.5B LLM (Fig. 5(b), more details in Appendix G). This phenomenon is a multi-task extension of the “Learning at Criticality” (LaC) mechanism, first identified in a single-task context (Cai et al., 2025). The LaC framework models this process as a continuous phase transition (Stanley & Ahlers, 1972; Wilson, 1983) where the policy collapses from exploring diverse reasoning paths to exploiting a single optimal one. Crucially, analysis revealed that near this critical point, the system’s exploratory power is maximized, and the reasoning path lengths become scale-invariant—a hallmark of continuous phase transitions. Beyond this transition, the policy collapses onto fixed optimal paths; the associated generalization drop is possibly linked to overfitting within the RLVR process (Dong et al., 2025; Gui & Cheng, 2025; Shypula et al., 2025; Zheng et al., 2025). This context explains the smoother gradient in the LLM. Compared to the step-like jumps in CoNet, the LLM’s learning curve is a classic manifestation of finite-size effects (Fisher & Barber, 1972): it has a more constrained set of distinct reasoning paths for any given task, which rounds out the idealized sharp transition into a steep but continuous curve.
The puzzle is why this delicate phenomenon survives in a chaotic multi-task setting. Our sparse-web model explains this through topological isolation: a node on the expanding frontier of a $\left<k\right>\approx 2$ network is statistically isolated, decoupling it from system-wide competition. Shielded from interference, a new skill can be efficiently integrated via a clean phase transition. This perspective also provides a microscopic explanation for aggregate policy collapse in the standard RLVR (Chen et al., 2025; Yue et al., 2025; Cui et al., 2025). We hypothesize this is not a monolithic failure but the statistical sum of asynchronous, microscopic collapses at the frontier, a hypothesis supported by the LaC finding that training just beyond criticality risks overfitting-induced collapse.
5 Annealed-RLVR: An Algorithm Born from the Sparse-Web Theory
(a) Training Sample (512 Problems)
(b) Minerva
(c) AIME 2024/2025
Figure 6: Annealed-RLVR Outperforms Standard RLVR. This figure compares the best@k accuracy curves of our Annealed-RLVR ("Annealed") with the standard RLVR baseline ("RLVR"). The evaluation is performed on (a) an in-distribution set (Randomly Selected 512 Training Problems) and two out-of-distribution (OOD) sets: (b) the Minerva dataset and (c) the AIME 2024/2025 datasets. The results show that Annealed-RLVR consistently achieves superior performance across all in-distribution and OOD benchmarks, confirming that the intervention fosters improved generalization.
Our sparse-web theory reveals a core narrative of RLVR training: the system drives itself into a maximally frustrated state that is both a perilous bottleneck and a moment of peak exploratory potential (Fig. 6(b), blue curve). This insight motivates Annealed-RLVR, a theory-driven intervention that leverages the high-entropy frustrated state to guide the model toward a more globally optimal configuration (Cheng et al., 2025; Chen et al., 2025).
The algorithm operationalizes this concept with a precisely timed intervention. This intervention is triggered by monitoring the macroscopic signatures of the maximally frustrated state, which, as shown in Fig. 1, correspond to the crossover regime characterized by the “knee” of the reward curve (when accuracy gains plateau) and the bottom of the V-shaped response length curve. At this juncture, a brief SFT “heating” phase begins. The timing is critical: because the “skill islands” have not yet solidified into a single web, the system is uniquely robust to a targeted shock that would otherwise cause catastrophic forgetting—a theoretical claim confirmed by our experiments (Fig. 4). The SFT thus functions not as a disruptive overwrite but as a corrective jolt to break the policy’s premature commitment to suboptimal routes. We implement this by surgically applying SFT only to problems with low accuracy (e.g., <0.1<0.1) for which correct solutions exist. Immediately following this targeted heating, the “cooling” stage resumes standard RLVR, allowing the now more pliable policy to re-settle into a more robust and integrated final state.
We validated this approach on the DeepSeek-R1-Distill-Qwen-1.5B model, trained on the DeepScaleR-Preview-Dataset . Our method, Annealed-RLVR, was compared against the standard RLVR baseline where the policy is optimized using a variant of the Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024; Liu et al., 2025; Yu et al., 2025). The results (Fig. 6) confirm our theory. Best@k accuracy curves demonstrate that Annealed-RLVR outperforms standard RLVR on both in-distribution evaluation set (512 randomly selected training problems) and on out-of-distribution (OOD) benchmarks, including the Minerva and AIME 2024/2025 dataset. The histograms (see Fig. 13) reveal the underlying mechanism: the SFT “heating” provides a crucial performance boost to the most difficult problems, lifting their success rates from near-zero and making them “visible” to subsequent RLVR exploration. The RLVR “cooling” phase then efficiently guides these newly accessible problems toward mastery. This results in a systematic reduction of unsolved problems and a corresponding population shift towards solved ones, confirming that our intervention guides the model past the inherent bottlenecks of standard RLVR to achieve superior reasoning. More details are given in Appendix H and I.
6 Conclusion and Outlook
We have introduced a sparse-concept-net theory to explain the puzzling behaviors of the RLVR training, including the two-stage learning curve, V-shaped response curves, and vulnerability to catastrophic forgetting. Our theory proposes that these diverse phenomena are consequences of one underlying process: the formation of a sparse concept web, with the averaged degree pinned close to two. This physical model directly explains the V-shaped curve as a shift from local optimization to global integration and recasts catastrophic forgetting as the severing of the web’s critical, bridge-like connections.
A natural consequence of this sparse structure is a fundamental RLVR mechanism for forgetting and learning within the microscopic training process. We find that the degree-2 constraint first forces the system into a maximally frustrated state at the onset of slow learning, a phase dominated by frustration-induced forgetting as skills are competitively suppressed. Subsequently, the same sparse structure enables a cleaner growth mechanism: phase-transition-like learning, where new knowledge is efficiently integrated at the web’s expanding frontier , yet this very per-skill collapse onto a single high-reward path also accumulates over time into a macroscopic policy collapse and a global loss of solution diversity when training proceeds too far beyond criticality.
This framework is not just descriptive; it is prescriptive. It identifies the maximally frustrated state—the peak of competitive forgetting—as the ideal moment for SFT intervention. Our Annealed-RLVR algorithm is born from this insight. By applying a timed, corrective SFT shock, Annealed-RLVR mitigates the policy collapse and, as our experiments confirm, unlocks a superior final reasoning capability. Ultimately, this work provides a new lens for understanding RLVR: not as black-box optimization, but as a predictable process of structural self-organization, paving the way for more principled methods for engineering the advanced reasoning capabilities of future AI.
Looking forward, a central long-term goal is the empirical mapping of the complete, microscopic reasoning graph from large-scale foundation models. This remains a significant undertaking, requiring substantial computational resources and novel algorithms for semantic chunking and clustering. The payoff for this effort, however, would be transformative, producing a powerful new analytical object: a detailed map of an AI’s internal world, composed of the reasoning graph and its emergent concept web. While the immediate application of this map would be to empirically test our sparse-web hypothesis, its long-term impact is far broader: charting this internal world is a critical step toward genuinely explainable LLMs and the robust safety of future artificial intelligence.
7 Related Work
Our work intersects with three domains: training paradigms for LLM reasoning, the modeling of reasoning as a complex network, and the study of RLVR’s failure modes. While building on the standard RLVR paradigm (Guo et al., 2025; Team et al., 2025; Yang et al., 2025), our goal is to provide a theoretical-physics-motivated theory for its perplexing macroscopic dynamics. To this end, we engage with the view of reasoning as a complex network; however, our approach differs from prior work that empirically extracts static reasoning graphs post-hoc (Xiang et al., 2025; Cabannes et al., 2024; Minegishi et al., 2025). Inspired by the core logic of the RG, we bypass the intractable graph-construction challenge and instead hypothesize a simple organizational principle for the emergent, coarse-grained concept web with $\left<k\right>\approx 2$, enabling us to study the universal behaviros of RLVR. This sparse-web perspective provides a unified origin for seemingly distinct failure modes like catastrophic forgetting (Li & Hoiem, 2017; Ding & Wang, 2025) and policy collapse (Yue et al., 2025; Cui et al., 2025), framing the former as the severing of the network’s fragile bridges, and the latter as a consequence of a phase-transition-like learning dynamics at its frontier (Power et al., 2022; Lin et al., 2025; Cai et al., 2025). An important line of research has made valuable progress on these failures using algorithmic solutions (Chen et al., 2025). Our work contributes a complementary, theory-driven perspective: we leverage the sparse-web model to design Annealed-RLVR, an intervention that targets the problem’s underlying structural root. See Appendix B for extended related work.
8 Reproducibility Statement
To ensure full reproducibility, we provide a complete experimental blueprint in the Appendix: Sec. I details the LLM training pipeline—models, hyperparameters, and SFT dataset construction, while Sec. J describes our evaluation protocol. Our code repository of CoNet can be available at: https://anonymous.4open.science/r/CoNet-83A4. The repository includes a README.md file that details the functionality of each key Python script. The training process is described in Appendix D and Appendix H.
Appendix A The Use of Large Language Models (LLMs)
This work utilized Large Language Models (LLMs) to assist in the preparation of the manuscript, primarily for improving language clarity and readability. For the accompanying CoNet repository, we also employed LLMs to generate Python codes, such as the training algorithm, measuring network properties and visualizing the concept web and its largest connected component. All LLM-generated outputs, including text and code, were meticulously reviewed, edited, and verified by the authors to ensure their accuracy and suitability for this work.
Appendix B Extended Related Work
System-2 Reasoning in LLMs. The dual-process theory, distinguishing between fast, intuitive System-1 thinking and slow, deliberate System-2 reasoning (Kahneman, 2011), has become a valuable framework for analyzing LLM behavior. While standard LLMs often exhibit System-1-like responses, significant research has focused on eliciting deliberate System-2 reasoning. Prompting strategies such as Chain-of-Thought (CoT) (Wei et al., 2022) and its variants like Self-Consistency (Wang et al., 2022) and Tree of Thoughts (ToT) (Yao et al., 2023) encourage models to generate explicit reasoning steps. Other works aim to build this deliberative capacity into the model’s architecture or training process through self-correction and reflection (Shinn et al., 2023; Madaan et al., 2023). Our work complements these efforts by shifting the focus from eliciting reasoning to understanding the emergent mechanism of how this capability develops internally during training.
Complex Network Analysis of Neural Models. Applying network science to understand neural networks is a long-standing idea, previously used to study the topology and robustness of smaller models (Simard et al., 2005; Estrada & Hatano, 2008). With the advent of Transformers, researchers have begun analyzing attention patterns as graphs (Vig, 2019) to better understand information flow (Abnar & Zuidema, 2020; Chefer et al., 2021). More recently, studies have explored the topological properties of neural networks, identifying optimal graph structures (You et al., 2020) and the formation of essential sparse subnetworks (Frankle & Carbin, 2019; Filan et al., 2021). Complementing these static analyses, recent work has modeled the training process itself as the evolution of complex networks (Danovski et al., 2024; Rocchetti, 2025). A recent study by (Wang et al., 2024; Minegishi et al., 2025; Matsutani et al., 2025) shares our goal of analyzing the network structures that underpin System-2 reasoning, though they do so by empirically extracting a graph from hidden-state representations. In contrast, our CoNet framework utilizes a fixed, abstract concept space, which allows us to specifically isolate and analyze the dynamic mechanism of how this reasoning structure emerges during training.
Appendix C The Reasoning Graph: From Empirical Extraction to a Unified Concept Web
The idea of a reasoning graph has emerged in the literature as a powerful framework for understanding the chain-of-thought process in LLMs for a given task. This graph is not explicitly encoded in the Transformer’s static parameters but is an emergent property of its inference dynamics. The standard model for its construction is based on a key observation: an LLM’s generative process is guided by a few decision points of high Shannon entropy, or “forking tokens” (Fig. 7). This allows for a deconstruction of a chain-of-thought: stable, low-entropy text chunks form the graph’s nodes (concepts), while the probabilistic choices at these high-entropy forks form the weighted, directed edges.
Figure 7: Illustration of the LLM reasoning process. A chain-of-thought is composed of stable, low-entropy text chunks (concepts), connected by high-entropy ’forking tokens’ that represent decision points. The LLM reasoning process can be deconstructed into stable “concepts” (e.g., step 3) linked by high-entropy decision points. The probabilities at these junctures (e.g., “next” vs. “tens”) represent transition weights on an underlying reasoning graph, a view supported by convergent findings on the role of high-entropy tokens. This figure is adapted from Ref. Cai et al. (2025).
Recent work, notably by (Minegishi et al., 2025), has developed detailed methods to empirically construct these task-specific reasoning graphs. Their process involves generating a large corpus of reasoning chains, segmenting them into individual steps, and then computing a vector representation for each step by averaging the hidden states of its tokens at a specific Transformer layer. Finally, a clustering algorithm (e.g., K-means) is used on the full set of step vectors, with the resulting cluster centroids being defined as the graph’s nodes. By analyzing graphs constructed this way, they found that more advanced models exhibit distinct topological properties, such as larger diameters, which correlate with improved performance. However, the cluster centroids derived from K-means offer limited interpretability as stable reasoning concepts, making their correspondence to the model’s internal processes unclear. Moreover, the induced graph’s structure is unstable, varying significantly with the choice of model, dataset, segmentation, and representation layer. Hence, this graph cannot track evolving representational dynamics during RLVR training process. How to construct a principled model of an LLM’s internal reasoning process remains a formidable open problem.
While analyzing these static, task-specific graphs is useful, it leads to a deeper question: what is the underlying, universal knowledge structure that allows an LLM to generate so many different reasoning graphs across a multitude of tasks? It follows that these empirically observed reasoning graphs are best understood as small projections of a much vaster, latent concept web. This web represents the entirety of the model’s abstract knowledge, and the ultimate goal is to understand its structure.
Figure 8: From Intractable LLM Dynamics to a Minimal CoNet Model. The reasoning process in an LLM (left) can be viewed as a traversal on a high-dimensional latent graph, where path probabilities are determined by the complex, autoregressive Transformer. Directly analyzing this graph’s evolution is intractable. CoNet (right) provides a minimal abstraction by replacing the intractable latent graph with a fixed K-regular random graph and simplifying the generation policy to a learnable Markov chain. This creates a tractable “computational microscope” to precisely analyze how the transition policy evolves under reinforcement, isolating the core dynamics of reasoning path formation. This figure is adapted from Ref. Cai et al. (2025).
However, attempting to reverse-engineer this universal concept web from its sparse, task-specific projections presents formidable obstacles. The difficulties are multifold. First, there is an issue of scale: the true concept web is of astronomical size, and any extracted graph is a small sample. Second, the mapping itself is ambiguous: the mapping from a low-level token sequence to a high-level semantic concept is inherently many-to-one. Third, this already-ambiguous mapping is itself dynamic, as it can shift during RLVR training. These challenges collectively lead to the observable node stability problem in practice, making it nearly impossible with current methods to track the web’s true topological evolution.
It is precisely this observational challenge that our paper bypasses with a two-pronged, indirect approach. First, we form a theoretical hypothesis about the consequences of a specific topological structure (i.e., sparsity). Second, we use a minimal model (CoNet) that captures the essential physics of the system to study these consequences in a controlled environment. Thus, while the direct construction of the concept web remains an open problem, our work proceeds from a central hypothesis about its fundamental structure. This hypothesis—that the web is profoundly sparse with an average degree near two—serves as the theoretical foundation for our paper. From it, we first demonstrate how this assumption of sparsity explains the known macroscopic dynamics of RLVR. We then use it to deduce previously overlooked microscopic mechanisms, namely frustration-induced forgetting and phase-transition-like learning. Finally, armed with these new insights, we explore how this deeper understanding can be leveraged to design a novel, improved RLVR algorithm.
Appendix D CoNet: A Minimal Model for RLVR dynamics
To bypass the intractable node stability problem, we leverage the Concept Network Model (CoNet). CoNet was originally introduced in (Cai et al., 2025) to study the dynamics of single-task RLVR, but its core design makes it perfectly suited for our investigation. By positing a fixed space of concept nodes (a static graph structure) and isolating all learning dynamics to the evolving policy that navigates it, CoNet allows us to sidestep the challenges of direct graph extraction and precisely measure how the reinforcement signal sculpts reasoning paths. The model represents the LLM’s concept space as an abstract K-regular random graph, with the reasoning process simplified to a policy-guided Markovian random walk, as shown in Fig. 8. The transition policy, $\pi_{\theta}(j|i)\propto\theta_{ij}$, is governed by a set of learnable parameters $\theta_{ij}$ that represent the strength of the connection from node i to node j. This policy is updated via RLVR on successful paths from a question (Q) to an answer (A).
This design offers a clear cognitive analogy. A single-step hop on the graph represents “System 1” thinking—fast, intuitive associations. A multi-step traversal, guided by the learned policy, represents “System 2” reasoning—a deliberate, compositional chain of thought. It is crucial to note that CoNet is not a model of the Transformer’s parameters themselves; rather, it models the emergent semantic network induced by inference—the structure that governs how questions are connected to answers through chains of concepts.
CoNet was originally introduced to model the phenomenon of Learning at Criticality (LaC), which was identified in single-task RLVR training (Cai et al., 2025). The central finding of LaC is that for a model trained on extremely sparse data (e.g., a single exemplar), generalization performance on unseen problems peaks precisely at a sharp, sigmoidal learning transition, before declining due to overfitting. The original work used CoNet to establish that this dynamic is a learning phase transition, providing a physical model for this peak in generalization.
The central methodological advance of this paper is to deploy this existing model in a richer, multi-task setting with numerous concurrent Q-A pairs. This extension is crucial: it transforms CoNet from a model of isolated skill acquisition into an ideal theoretical laboratory for studying skill integration. In this regime, the model is compelled by competing reinforcement signals to discover a shared, reusable structure. As Fig. 1 shows, this multi-task CoNet stunningly reproduces the key macroscopic signatures of the 1.5B LLM. This correspondence validates our use of CoNet as a minimal model for exploring the network-level mechanisms of frustration and web formation.
Appendix E CoNet Implementation Details
The policy was optimized using a variant of the Group Relative Policy Optimization (GRPO) algorithm ((Shao et al., 2024; Liu et al., 2025; Yu et al., 2025)). For a given Q-A pair, $n_{\text{rollout}}=128$ reasoning paths (indexed by $m$) were sampled, with each path’s exploration capped at a maximum length of 20 steps. Each path received a reward based on its success, and its advantage $A_{m}$ (relative to the average reward over all 128 rollouts) was used to guide the update of the policy parameters $\theta_{ij}$. The update rule is given by:
Appendix F Structural Evolution of the Concept Web in CoNet
This section provides direct, quantitative evidence from the CoNet model for the two-stage learning dynamic—a fast acquisition process followed by a slow integration phase—that underlies our theory. Fig. 9 serves as the central exhibit for this analysis, tracking two key structural properties over the course of training: the total number of disconnected solution clusters (the “skill islands”) and the size of the single largest cluster, namely the concept web.
Figure 9: The Formation of the Concept Web in CoNet. This figure illustrates the evolution of network clusters during training, showing a structural reorganization from isolated skills to a unified conceptual web. In the initial phase, the Cluster Number (orange curve) spikes, indicating the rapid discovery of numerous disconnected “skill islands”. The peak marks a critical structural juncture. Following this peak, the Cluster Number steadily declines as these islands begin to coalesce. This merging process is confirmed by the corresponding monotonic growth in the Max Cluster Size (blue curve), which represents the formation of a single giant component. By the end of training, the network is dominated by this large, interconnected structure.
The slow-learning phase is thus defined by the coalescence of these islands into a unified concept web. The plot reveals the two complementary trends that define this stage: a steady decline in the total number of clusters (orange curve) and a simultaneous, relentless growth in the size of the largest cluster (blue curve). The causal link is direct: the largest component grows precisely by absorbing smaller, independent clusters, which necessarily reduces the total cluster count. This structural evolution culminates in a state where a single, giant component—the functional concept web—dominates the network. This process provides a direct, mechanical explanation for the V-shaped response length signature: Stage 1 shortens paths by finding local solutions, while Stage 2 lengthens them by forcing the policy to traverse longer, connective paths between previously distant concepts.

### v3

@@ Appendix E Structural Evolution of the Concept Web in CoNet
Figure 11: Direct visualization of inverse-tree freezing in CoNet. Network snapshots at the three training steps marked on Fig. 10, retaining only edges with transition probability $p>0.95$. Green = question nodes, red = answer nodes, blue = intermediate nodes; edge colour encodes the transition probability (right colour bar). Unlike Fig. 4 of the main text, which extracts only the largest inverse tree at each step, this figure shows all surviving high-probability edges, exposing the full landscape of connected components throughout training. (a) Step 20: multiple short, disjoint subtrees nucleate independently across the graph—the nascent “skill islands”—coinciding with the peak of the cluster-number curve in Fig. 10 and the onset of the maximally frustrated state. (b) Step 50: a dominant component has emerged through the merging of compatible paths at shared nodes, while a handful of small subtrees have yet to be absorbed; this is the regime in which frustration-induced forgetting is most active. (c) Step 800: the surviving high-probability edges have coalesced into a single connected inverse tree with average degree $\langle k\rangle\approx 2$, the frozen-in concept web that supports the trained policy.
These three macroscopic indicators (Fig. 10) are made microscopically concrete by the network snapshots in Fig. 11, which visualize the same three training steps directly. The Step 20 snapshot makes the “proliferation of skill islands” literal: many short, disjoint paths populate the graph in parallel, exactly as expected when nucleation dominates and merging has not yet set in. The Step 50 snapshot shows the structural fingerprint of the maximally frustrated state: a dominant connected component has begun to grow by welding compatible paths at shared nodes, but several subordinate islands persist, sustaining the inter-task competition that drives frustration-induced forgetting. By Step 800, the high-probability edge set has frozen into a single, sparse inverse tree—the concept web—whose multi-input, single-output topology and average degree $\langle k\rangle\approx 2$ are precisely the structural ingredients that the main text identifies as responsible for both the V-shaped response-length signature and the bridge-node fragility under SFT.
Appendix F Geometric Origin of the Rising Branch of Response Length
The main text argues that the lengthening of correct responses during Stage 2 is a direct geometric consequence of the inverse tree: with average degree $\langle k\rangle\approx 2$ and out-degree close to $1$ at almost every node, the network has no short-range shortcuts, so as the inverse tree extends to cover more question–answer pairs the typical question-to-answer geodesic grows monotonically. This appendix provides the corresponding microscopic evidence in CoNet: tracking the largest inverse tree alongside the distribution of correct path lengths during Stage 2 shows that the two evolve together, exactly as the inverse-tree picture predicts.
(a) Correct-response length distribution
@@ Appendix F Geometric Origin of the Rising Branch of Response Length
Figure 12: Inverse-tree growth and the rising branch of response length in CoNet. (b, c) Largest connected component at $p>0.95$ at Step 50 and Step 800, drawn with the same colour and marker scheme as Fig. 11. The inverse tree expands substantially between the two steps—absorbing additional question–answer pairs through merging at shared nodes—while keeping its sparse, multi-input/single-output character: average degree stays $\langle k\rangle\approx 2$ (cf. Fig. 10). (a) Distribution of correct solution path lengths over the same training window (raw counts, with the probability density shown in the inset). As the inverse tree grows, the length distribution shifts decisively to the right; both the mean and the upper tail move outward, while the support remains bounded above by the maximum episode length.
The rightward shift of the length distribution in Fig. 12a is not an incidental feature of CoNet’s hyperparameters but the direct macroscopic signature of the geometric constraint imposed by the frozen inverse tree. Once Stage 2 begins, new question–answer pairs are no longer mastered by discovering fresh short-cut paths in the underlying random graph; they are mastered by being grafted, through merging, onto the existing inverse tree at one of its branching nodes. Each such grafting attaches an additional segment of length set by the graph distance from the new question node to the nearest node already on the tree, and this distance is itself bounded below by the tree’s growing radius. As a result, the typical solution path length must increase together with the size of the inverse tree, exactly as observed when the largest component expands from Step 50 (Fig. 12b) to Step 800 (Fig. 12c).
This sparse-tree-geodesic argument also clarifies what the rising branch is not. It is not driven by the model adopting longer reasoning styles, nor by a switch to harder problems, nor by stochastic noise on already-mastered paths: in CoNet none of these can vary, since the graph, the problem set, and the policy parameterisation are all fixed by construction. The only quantity that changes between Step 50 and Step 800 is the topology of the high-probability sub-network, and yet the response-length histogram shifts decisively. The rising branch of the V is therefore a structural readout of inverse-tree growth, in line with the corresponding mechanism stated in the main text. Self-correction or reflection-style cycles are fully compatible with this picture: they are sparse, local loops on the same backbone, occasional rather than dense, and do not appreciably change the global average degree from $\langle k\rangle\approx 2$.